# GRAPHITE: Graph-Based Interpretable Tissue Examination for Enhanced Explainability in Breast Cancer Histopathology

### v1

@@ 3.9 Clinical Utility and Net Benefit Analysis
The clinical utility of the evaluated XAI methods was further assessed using NBC, with the AUDC metric serving as a quantitative measure of clinical effectiveness. GRAPHITE-V2 achieved the highest AUDC of 4.17e+5, indicating the highest net benefit across decision thresholds (Figure 9). This result highlights the potential of GRAPHITE-V2 to improve clinical decision-making by balancing true positive predictions with a minimised cost of false positives. GRAPHITE-V1 and GRAPHITE-Base also show high AUDC values, further validating the robustness of the GRAPHITE framework. Traditional methods such as GradCAM and FullGrad demonstrated lower AUDC scores, reflecting their limited effectiveness in scenarios where threshold-dependent decisions are critical.
@@
Figure 8: Comparative visualisation of breast cancer TMA cores using GRAPHITE, attention-based methods and Grad-CAM, alongside the actual TMA core and pathologist-annotated mask. Each row represents a different TMA core. The GRAPHITE method demonstrates improved alignment with annotated masks compared to other methods, capturing more relevant regions of interest and enhancing interpretability in breast cancer analysis.
Figure 9: Net Benefit Curves for various XAI methods, illustrating the clinical utility of each method across a range of decision thresholds. The net benefit is plotted as a function of threshold, showing the trade-off between true positives and the cost of false positives for each method. The AUDC is provided in the legend for each method as a quantitative measure of the model’s utility. GRAPHITE-V2 achieves the highest AUDC 4.17e+5), followed by GRAPHITE-V1 (4.08e+5) and GRAPHITE-Base (4.15e+5), indicating superior net benefit across thresholds. These results highlight the advantage of the GRAPHITE variants over other XAI methods, such as Attention (AUDC = 4.05e+5) and FullGrad (AUDC = 3.56e+5). The dashed line represents the “no net benefit” baseline, where a model yields no additional benefit over treating all or none of the cases. The higher net benefit values for GRAPHITE variants underscore their potential effectiveness in clinical decision-making for breast cancer analysis.
@@ 4 Discussion
When compared to existing XAI methods, GRAPHITE’s advantages become particularly apparent. Traditional approaches like GradCAM (AUROC = 0.86) and attention-based methods (AUROC = 0.92) struggle to provide consistent interpretations across different scales and tissue types. GRAPHITE’s hierarchical approach and adaptive fusion mechanism address these limitations, resulting in more reliable and clinically meaningful visualisations. The framework’s superior performance in both discriminative power and interpretability metrics suggests that it successfully bridges the gap between model accuracy and clinical utility, a crucial consideration for real-world applications.

### v2

@@ 3.9 Clinical Utility and Net Benefit Analysis
The clinical utility of the evaluated XAI methods was further assessed using NBC, with the AUDC metric serving as a quantitative measure of clinical effectiveness. GRAPHITE-V2 achieved the highest AUDC of 4.17e+5, indicating the highest net benefit across decision thresholds (Figure 10). This result highlights the potential of GRAPHITE-V2 to improve clinical decision-making by balancing true positive predictions with a minimised cost of false positives. GRAPHITE-V1 and GRAPHITE-Base also show high AUDC values, further validating the robustness of the GRAPHITE framework. Traditional methods such as GradCAM and FullGrad demonstrated lower AUDC scores, while the model-agnostic methods showed considerably less utility in this clinical context. LIME achieved an AUDC of 2.70e+5, and SHAP performed poorest with an AUDC of 1.72e+5. These low values for LIME and especially SHAP align with their performance in other metrics and underscore their limited effectiveness in guiding threshold-dependent clinical decisions for this specific task compared to the GRAPHITE variants.
@@ 4 Discussion
When compared to existing XAI methods, GRAPHITE’s advantages become particularly apparent. Traditional approaches like GradCAM (AUROC = 0.86) and attention-based methods (AUROC = 0.92) can struggle to provide consistent interpretations across different scales and tissue types. Similarly, widely used model-agnostic methods such as LIME and SHAP—though flexible—demonstrated limitations in our evaluation (e.g., SHAP AUROC = 0.61), likely due to challenges in capturing complex spatial dependencies within intricate tissue patterns. GRAPHITE’s hierarchical approach and adaptive fusion mechanism appear better suited to address these histopathology-specific challenges, resulting in more reliable and clinically meaningful visualisations. The framework’s superior performance across discriminative power, interpretability metrics, and clinical utility assessments (AUDC) compared to this broad range of techniques suggests that it successfully bridges the gap between model accuracy and clinical utility, a crucial consideration for real-world applications.